OpenStudy (anonymous):

MEDAL AND fan!!!!!!!!!!! Write the equation of the line that passes through (3, 4) and (2, –1) in slope-intercept form

OpenStudy (jdoe0001):

\(\bf \begin{array}{lllll} &x_1&y_1&x_2&y_2\\ % (a,b) &({\color{red}{ 3}}\quad ,&{\color{blue}{ 4}})\quad % (c,d) &({\color{red}{ 2}}\quad ,&{\color{blue}{ -1}}) \end{array} \\\quad \\ % slope = m slope = {\color{green}{ m}}= \cfrac{rise}{run} \implies \cfrac{{\color{blue}{ y_2}}-{\color{blue}{ y_1}}}{{\color{red}{ x_2}}-{\color{red}{ x_1}}} \\ \quad \\ % point-slope intercept y-{\color{blue}{ y_1}}={\color{green}{ m}}(x-{\color{red}{ x_1}})\qquad \textit{plug in the values and solve for "y"}\\ \qquad \uparrow\\ \textit{point-slope form}\)

OpenStudy (anonymous):

slope is 5

OpenStudy (jdoe0001):

yeap so use the slope of 5 and plug it in the point-slope form using the \(x_1\ and\ y_1\) values alongside

OpenStudy (anonymous):

the y int is -9

OpenStudy (jdoe0001):

\(\bf \begin{array}{lllll} &x_1&y_1&x_2&y_2\\ % (a,b) &({\color{red}{ 3}}\quad ,&{\color{blue}{ 4}})\quad % (c,d) &({\color{red}{ 2}}\quad ,&{\color{blue}{ -1}}) \end{array} \\\quad \\ % slope = m slope = {\color{green}{ m}}= 5 \\ \quad \\ % point-slope intercept y-{\color{blue}{ 4}}={\color{green}{ 5}}(x-{\color{red}{ 3}})\qquad \textit{plug in the values and solve for "y"}\\ \qquad \uparrow\\ \textit{point-slope form}\)

OpenStudy (anonymous):

-11

OpenStudy (anonymous):

y= 5x-11 (slope intercept form)

OpenStudy (anonymous):

answer should be y=5x-11

OpenStudy (anonymous):

Write the equation of the line that passes through (–1, 5) and has a slope of 3 in point-slope form

OpenStudy (anonymous):

is it y-5=3(x+1)

OpenStudy (anonymous):

follow @jdoe0001